An angle's measure is 24 degrees less than that of its supplement. how would I measure find the measure of the triangle
Answer by Earlsdon(6294)

Well, let the angle be A. The supplement of angle A is (180 - A) degrees. The problem says that the measure of angle A is 24 degrees less than (180-A).
So we have: A = (180-A) - 24. Solve for A
A = 180 - A - 24 Add A to both sides.
2A = 180 - 24
2A = 156 Divide both sides by 2.
A = 78 degrees.
A = (180-78)-24
A = (102)-24
A = 78 degrees.
Check:
A = (180-A)-24
A = (180-78)-24
A = (102)-24
A = 78